# It’s Not Easy Being Green: On the Energy Efficiency of Programming Languages

### v3

@@ IV-B Programming Languages and Implementations
Figure 5: Programming Languages can exhibit different performance characteristics depending on the implementation used (§IV-B). Here, we compare interpreters (CPython, Lua) to their JIT equivalent (PyPy, LuaJIT) on the same CLBG benchmark sources. We normalize to the interpreter results. We find around 5×5$\times$5 × and 1.25×1.25$\times$1.25 × decreased execution time and energy consumption on average for LuaJIT over Lua and PyPy over CPython, respectively. These results demonstrate the impact of choice of programming language implementation versus the language itself.
@@
Figure 5 compares languages across multiple implementations on the same set of benchmarks, namely CPython/PyPy and Lua/LuaJIT, and shows that the same language can have vastly different performance characteristics, depending on the implementation used. On average for these benchmarks, PyPy is about 1.25×1.25$\times$1.25 × faster than CPython, and LuaJIT is 5×5$\times$5 × faster than the Lua interpreter.
Not only can programmers choose a specific implementation over another, implementation specific properties can cause significant differences in performance and energy consumption, and should be taken into account in a fair comparison. We discuss two of those properties below, just-in-time warmup and garbage collection.
IV-B1 Just-in-Time (JIT) Warmup
JIT program execution is typically split into a startup phase and a delayed steady state phase of peak performance. As a JIT program executes, it determines which code paths are frequently executed, and compiles them into machine code. As the program continues to revisit these hotspots, performance of further iterations improves. This system allows for a good startup time to peak performance tradeoff. However, existing research has already shown that a steady state of peak performance is not always reached [24, 23]. In fact, one of these studies [24] uses the same CLBG benchmark suite and demonstrates that for some benchmarks, a steady state is never reached.
(a) mandelbrot
(b) fannkuch-redux
Figure 6: Measuring just the first iteration is not indicative of Java’s (OpenJDK) performance for long-lived applications (§IV-B1). We gradually increase the number of iterations, observing decreased per-iteration execution time on several benchmarks, including (a) mandelbrot and (b) fannkuch-redux shown above. This effect could be due to startup and cleanup time and/or a JIT warmup effect. The curve flattens as we average over more iterations.
Factors other than JIT warmup may also affect the first iteration of short-lived benchmarks, such as bytecode interpretation or memory and garbage collection initialization. To verify this, we wrap Java benchmarks in a loop and gradually increase the number of in-process iterations. Figure 6 shows that for the mandelbrot and fannkuch-redux benchmarks, the first iteration’s execution time is respectively 2.7×2.7$\times$2.7 × and 1.4×1.4$\times$1.4 × slower than when averaging over 15 iterations. Generally, we see that for these and other benchmarks, averaging over ten or more iterations amortizes the first iteration’s impact: across the CLBG benchmarks, we observe a mean of 80% improved energy efficiency and execution time.
IV-B2 Garbage Collection
An important property of higher-level languages and their implementations is garbage collection, which imposes significant performance overhead [25]. Out of the 13 language implementations discussed in this paper, 10 use a garbage collector. Garbage collectors can typically be configured and tuned for each application to offer maximal performance, which is standard practice in real-world applications to optimize for application throughput and/or latency. For example, Go’s garbage collector can be disabled entirely with the GOGC environment variable. When doing so, we observe a 2.8×2.8$\times$2.8 × speedup on the allocation-intensive binary-trees benchmark, and a 1.23×1.23$\times$1.23 × speedup on mandelbrot.
IV-C Decomposing Energy into Power and Time
@@
Fundamentally, energy consumption is the product of power and time. Figure 7 decomposes these two factors. To test if the choice of programming language implementation has an effect on power draw beyond execution time, we fix other factors that may affect power draw: we limit the program to a single core via Docker’s --cpuset-cpus argument, and pin CPUs to their minimum frequency with turbo mode disabled to ensure frequency scaling and throttling do not increase power draw variance in our measurements. Once these factors are controlled for, we measure runtime and energy consumption of each benchmark across all language implementations. These measurements yield equal power draw: $189.8\pm 0.5$ W. Hence, the impact of choice of programming language implementation (or programming language more generally) on energy consumption beyond execution time is negligible.
In theory, any language implementation with a runtime system running in parallel may affect energy consumption beyond variance in execution time, since increased core usage will increase power draw, as Section IV-D1 details below. This hidden increased parallelism could occur with a parallel garbage collector or just-in-time compilation threads executing alongside the main program. In practice, compilation threads are primarily active only during startup, and both are rarely significant compared to the main program’s execution and other general overhead costs. For this reason, we draw this edge in gray in our model.
IV-D Contributors to Power Draw
@@
(a) Average package (PKG) power draw by average number of active cores (§IV-D1). Log regression: $y=31\log_{2}x+246\left(R^{2}=0.72\right)$.
@@
(b) Average DRAM power draw by memory activity (§IV-D2), approximated using LLC misses. Linear regression: $y=1.68\cdot 10^{-8}x+12\left(R^{2}=0.93\right)$.
@@
Previous research has demonstrated that processor and memory energy usage are the main variable contributors to power draw [34]. We add these two factors to our model in Figure 8. For the studied programs on our experimental platform, the ratio of memory energy consumption to CPU energy consumption remains between 2 and 8%. This percentage is in line with the reported numbers in Pereira et al.
IV-D1 Number of Active Cores
Figure 9(a) shows the relationship between the number of active cores and power draw. Using multiple cores, energy usage grows up to around 2×2$\times$2 × the base power draw compared to using a single core. We fit a log curve to the data $\left(R^{2}=0.72\right)$, as the relationship does not appear linear, but results may vary across platforms. Denoting $\mbox{Power}\left(x\right)$ as the average power draw using $x$ cores, we find that, on our system, doubling the number of cores used increases average power draw by only roughly 31W:
@@ IV-D1 Number of Active Cores
In other words, parallelization overhead up to 87% is acceptable. On our experimental platform, aggressively parallelizing programs is nearly always an energy-efficient choice.
IV-D2 Memory Activity
Previous research has shown that roughly 40% of dynamic random access memory (DRAM) energy consumption is constant and required to refresh memory cells to keep the system running properly, while the remaining 60% is related to read and write activity [35]. Section III-B2 introduces last level cache (LLC) misses as a better proxy for memory activity: any read or write request that cannot be fulfilled by the LLC will have to go to DRAM. Figure 9(b) shows a strong linear correlation between LLC misses and increased power draw $\left(R^{2}=0.93\right)$.
On the other hand, memory activity is not something that is easily controlled by the programmer. While certain algorithms or data structures maximize cache locality, there is eventually a limit stemming from the LLC’s fixed, relatively small size. Programming languages or their implementations may impose additional overhead with a garbage collector, memory allocation overhead, or object memory layout. Because of the difficulty of controlling memory activity, as well as its comparatively smaller impact compared to CPU activity on power draw (2–8%), we draw this edge in gray in our model.
IV-E Application Implementations
Finally, we add the application implementations themselves, which have the most direct impact on energy consumption. Figure 1 presents the final causal model developed in this paper. Application implementation decisions dictate the level of parallelism, the algorithms used, and the data structures employed. We discuss two factors, concurrency and third-party library usage, below.
IV-E1 Concurrency
Section IV-D1 asserts that the number of cores used by an application is a major factor in energy consumption. Therefore, comparing benchmark implementations in different languages that use different numbers of cores is not a fair comparison. Level of parallelism imbalances are the main reason for the reported performance discrepancies between JavaScript and TypeScript. The TypeScript version of the mandelbrot benchmark takes 21×21$\times$21 × longer to run than its JavaScript implementation. Its execution is fully sequential, while the JavaScript version uses 28 cores on average.
@@ IV-E1 Concurrency
IV-E2 Third-Party Libraries
Choice of third-party libraries used has a significant impact on application performance. In fact, regular expression benchmarks are often poor candidates to make any comparison beyond the library used. This is the case for the CLBG regex-redux benchmark: it is 8.9×8.9$\times$8.9 × slower in its C++ implementation compared to the C version. This difference is entirely due to the choice of third-party library used: the C version uses the PCRE library, while the C++ version uses the Boost library. On this benchmark, Boost’s library performs significantly worse than PCRE. This outlier alone accounts for the entire reported gap between C and C++.
@@ IV-E2 Third-Party Libraries
Crossing Language Boundaries
Using third-party libraries also allows for easily crossing language boundaries, for instance to access lower-level system utilities or to obtain greater performance. In Python, using NumPy can yield up to 60,000×60,000$\times$60 , 000 × faster code when multiplying matrices [36, 22]. On a small micro-benchmark multiplying $000\times 10,$ 000 × 10 , 000 matrics, we find that NumPy is 84×84$\times$84 × faster than a naive three-loop C++ implementation, with $\times4$ lower energy consumption. Of course, this is because NumPy’s underlying matrix-multiply implementation is written in a low-level language and leverages many optimizations such as parallelism, vector instructions, blocking, or hardware-specific instructions. To some degree, using NumPy and other third-party libraries has become the “correct” way of writing Python code: leveraging low-level languages for compute-intensive tasks, while keeping Python’s simplicity. Even C++ may cross over to code originally written using C, Fortran, or direct Assembly for performance when using BLAS.
IV-F Other Factors
V Threats to Validity
@@
VI Related Work
Previous sections, notably Section II, discuss and critique the main line of research relating programming languages and energy consumption [9, 10, 11]. These studies discuss and rank programming languages based on execution time, energy consumption, and memory usage. Follow-on studies [38, 39] use external energy measurements rather than the on-chip measurements used here and in Pereira et al., but still suffer from the same limitations as the original work.
@@ VI Related Work
Previous research also addresses the energy consumption of different data structures in certain languages, notably Java collections [45, 46, 47], or Python data manipulation libraries [48]. Data structures can be interesting candidates to study through an energy-focused lens as they can have a substantial impact on locality and memory activity, which, as Section IV-D2 describes, are factors in energy consumption. However, programmers often have a limited choice of data structures that fit for their use case or algorithm.
Finally, past work has also investigated energy efficiency of concurrent programs, focusing on Haskell’s data sharing primitives [49] or Java’s thread management constructs [50], or analyzing the power to performance tradeoffs in lock-based synchronization [51]. As Section IV-D1 argues, these studies confirm that the power usage patterns of the processor and the machine as a whole are more complex with parallel execution.
VII Conclusion
@@
The software, benchmarks, and data that support this paper’s experiments are available on GitHub at github.com/nicovank/Energy-Languages.
@@ Acknowledgments
Emery Berger holds concurrent appointments as a Professor of Computer Science at the University of Massachusetts Amherst and as an Amazon Scholar. This paper describes work performed at the University of Massachusetts Amherst and is not associated with Amazon.

### v4

@@ VI-A RQ1: Do Some Language Implementations Consume More Energy Than Others?
Figure 7: Programming Languages can exhibit different performance characteristics depending on the implementation used (§IV-B). Here, we compare interpreters (CPython, Lua) to their JIT equivalent (PyPy, LuaJIT) on the same CLBG benchmark sources. We normalize to the interpreter results. We find around $5\times$ and $1.25\times$ decreased execution time and energy consumption on average for LuaJIT over Lua and PyPy over CPython, respectively.
According to our model, the only significant causal path from programming language to energy consumption is mediated by three factors: the language’s implementation, the application’s implementation, and execution time. Therefore, for identical application implementations, power draw variance can only be due to language implementation differences. Similarly, if the number of cores used is fixed, variance in energy consumption is uniquely determined by execution time.
@@
For the Same Language: It is well known that different implementations of the same language may have widely varying performance characteristics. Figure 7 compares languages across multiple implementations on the same set of benchmarks, namely CPython/PyPy and Lua/LuaJIT, and confirms that not only do the JIT implementations offer increased performance, they also provide corresponding energy savings. On average for these benchmarks, PyPy is about $1.25\times$ faster than CPython, and LuaJIT is $5\times$ faster than the Lua interpreter.
Note that different implementations of the same language, as well as different versions of the same implementation, may have different tooling, compatibility, or support. These differences do not affect our experiments. In our PyPy/CPython and Lua/LuaJIT evaluation, all benchmarks run successfully in both the interpreter and JIT implementation, with only one exception for each: pidigits does not work with PyPy due to a third-party library compilation failure, and fasta fails in LuaJIT due to a standard library utility rename.
@@
Across Many Languages and Implementations: To test if the choice of programming language implementation has an effect on power draw beyond execution time, we fix other factors that may affect power draw: we limit the program to a single core via Docker’s --cpuset-cpus argument, and pin CPUs to their minimum frequency with turbo mode disabled to ensure frequency scaling and throttling do not increase power draw variance in our measurements. Once these factors are controlled for, we measure execution time and energy consumption of each benchmark across all language implementations. These measurements yield equal power draw: $189.8\pm 0.5$ W. Hence, the impact of choice of programming language implementation (or programming language more generally) on energy consumption beyond execution time is negligible.
RQ1 Summary: Different language implementations can have widely varying performance characteristics, and by extension, energy efficiency. However, once accounting for external factors such as numbers of cores used, all benchmarks in all language implementations exhibit constant power draw. Hence, energy consumption is directly proportional to execution time. Faster language implementations will offer commensurate energy savings.
VI-B RQ2: What Are the Key Contributors to Power Draw Standing Between Choice of Programming Language and Energy Consumption?
@@
(a) Average package (PKG) power draw by average number of active cores. Regression: $y=31\log_{2}x+246\left(R^{2}=0.72\right)$.
@@
(b) Average DRAM power draw by memory activity, approximated using LLC misses. Regression: $y=1.68\cdot 10^{-8}x+12\left(R^{2}=0.93\right)$.
@@
Our causal model identifies number of active cores and memory activity as primary contributors to power draw. Section V details our improved measurement methodology, which allows us to accurately quantify to which extent these factors actually contribute to power draw.
Number of Active Cores: Figure 8(a) shows the relationship between the number of active cores and power draw. Using multiple cores, energy usage grows up to around $2\times$ the base power draw compared to using a single core. We fit a log curve to the data $\left(R^{2}=0.72\right)$, as the relationship does not appear linear, but results may vary across platforms. Denoting $\mbox{Power}\left(x\right)$ as the average power draw using $x$ cores, we find that, on our system, doubling the number of cores used increases average power draw by only roughly 31W:
@@
In other words, parallelization overhead up to 87% is acceptable. Therefore, on our experimental platform, aggressively parallelizing programs is nearly always an energy-efficient choice.
Memory Activity: Previous research has established that roughly 40% of dynamic random access memory (DRAM) energy consumption is constant and required to refresh memory cells to keep the system running properly, while the remaining 60% is related to read and write activity [36]. Section V-A introduces last level cache (LLC) misses as a better proxy for memory activity: any read or write request that cannot be fulfilled by the LLC will have to go to DRAM. Figure 8(b) shows a strong linear correlation between LLC misses and increased power draw $\left(R^{2}=0.93\right)$. However, this figure also highlights how much smaller the memory’s energy consumption is to the processor’s, with their ratio remaining from 2 to 8 % across all benchmarks.
RQ2 Summary: Number of active cores is the primary factor in increased power draw. Memory energy consumption is linearly related to memory usage, but significantly less important than processor energy consumption. Moreover, aggressive parallelization, even with high overheads, increases energy efficiency.
VI-C RQ3: Can Observed Anomalies in Prior Work Be Explained Through the Lens of Our Causal Model?
This section identifies four major sources of anomalies in prior work. For each anomaly, we indicate in brackets which causal path(s) in our model are responsible.
Concurrency [Application Implementation → {Number of Active Cores, Time}]: The number of cores used by an application is a major factor in energy consumption, as Section VI-B quantifies. Therefore, comparing benchmark implementations in different languages that use different numbers of cores is not a fair comparison. Level of parallelism imbalances are the main reason for the reported performance discrepancies between JavaScript and TypeScript. The TypeScript version of the mandelbrot benchmark is $21\times$ slower than its JavaScript implementation. Its execution is fully sequential, while the JavaScript version uses 28 cores on average.
@@ VI-C RQ3: Can Observed Anomalies in Prior Work Be Explained Through the Lens of Our Causal Model?
Third-Party Library Usage [Application Implementation → {Time, Number of Active Cores}]: Choice of third-party libraries used has a significant impact on application performance. In fact, regular expression benchmarks are often poor candidates to make any comparison beyond the library used. This is the case for the CLBG regex-redux benchmark: it is $8.9\times$ slower in its C++ implementation compared to the C version. This difference is entirely due to choice of third-party library: the C version uses the PCRE library, while the C++ version uses the Boost library. On this benchmark, Boost’s library performs significantly worse than PCRE. This outlier alone accounts for the entire reported gap between C and C++.
@@
Crossing Language Boundaries: Using third-party libraries also allows for easily crossing language boundaries, for instance to access lower-level system utilities or to obtain greater performance. In Python, using NumPy can yield up to $60,000\times$ faster code when multiplying matrices [37, 22]. On a small micro-benchmark multiplying $10,000\times 10,000$ matrices, we find that NumPy is $84\times$ faster than a naive three-loop C++ implementation, with $44\times$ lower energy consumption. Of course, this is because NumPy’s underlying matrix-multiply implementation is written in a low-level language and leverages many optimizations such as parallelism, vector instructions, blocking, or hardware-specific instructions. To some degree, using NumPy and other third-party libraries has become the “correct” way of writing Python code: leveraging low-level languages for compute-intensive tasks, while keeping Python’s simplicity. Even C++ may cross over to code originally written using C, Fortran, or direct Assembly for performance when using BLAS.
Just-in-Time (JIT) Warmup [Programming Language Implementation → Time]: JIT program execution is typically split into a startup phase and a delayed steady state phase of peak performance. As the program executes, the JIT runtime determines which code paths are frequently executed, and compiles them into machine code, eventually optimizing hotspots. However, existing research has already shown that a steady state of peak performance is not always reached [24, 23]. In fact, one of these studies [24] uses the same CLBG benchmark suite and demonstrates that for some benchmarks, a steady state is never reached.
(a) mandelbrot
(b) fannkuch-redux
Figure 9: Measuring just the first iteration is not indicative of Java’s (OpenJDK) performance for long-lived applications (§IV-E). We gradually increase the number of iterations, observing decreased per-iteration execution time on several benchmarks, including (a) mandelbrot and (b) fannkuch-redux.
Factors other than JIT warmup may also affect the first iteration of short-lived benchmarks, such as bytecode interpretation or memory and garbage collection initialization. To verify this, we wrap Java benchmarks in a loop and gradually increase the number of in-process iterations. Figure 9 shows that for the mandelbrot and fannkuch-redux benchmarks, the first iteration’s execution time is respectively $2.7\times$ and $1.4\times$ slower than when averaging over 15 iterations. Generally, we see that for these and other benchmarks, averaging over ten or more iterations amortizes the first iteration’s impact: across all CLBG benchmarks, we observe a mean of 80% improved energy efficiency and execution time.
Garbage Collection [Programming Language Implementation → {Time, Memory Activity}]: An important property of higher-level languages and their implementations is garbage collection, which imposes significant performance overhead [25]. Out of the 13 language implementations discussed in this paper, 10 use a garbage collector. Garbage collectors can typically be configured and tuned for each application to offer maximal performance, which is standard practice in real-world applications to optimize for application throughput and/or latency. For example, Go’s garbage collector can be disabled entirely with the GOGC environment variable. When doing so, we observe a $2.8\times$ speedup on the allocation-intensive binary-trees benchmark, and a $1.23\times$ speedup on mandelbrot. Disabling the garbage collector over all Go benchmarks yields an average of $1.15\times$ improved execution time and energy consumption.
RQ3 Summary: Variations in application implementations, notably parallelism and third-party library usage, explain observed anomalies between JavaScript/TypeScript and C/C++. Important language implementation specifics such as warmup and garbage collection must also adequately be accounted for: when doing so, the unexpectedly large gap between Go/Java and C significantly shrinks.
VII Threats to Validity
@@
VIII Other Related Work
Previous sections, notably Section II, discuss and critique the main line of research relating programming languages and energy consumption [9, 10, 11]. Follow-on studies [38, 39] use external energy measurements rather than on-chip measurements, but still suffer from the same limitations as the original work.
Georgiou et al. [40] explore the Energy Delay Product (EDP) of programming languages on Rosetta Code. The goal of EDP is to establish a trade-off between energy consumption and execution time. Our evaluation suggests that, in fact, optimizing for execution time also provides proportional energy consumption reduction. Georgiou et al.’s public data reveals that this conclusion is in line with their findings: even in the “optimize for energy” case, nearly all benchmark versions which are the fastest are also the most energy-efficient, with the rare exceptions attributable to short execution times, coarse measurement granularity, and/or missing multicore accounting. Our methodology leverages modern on-chip energy and performance counters, offering an analysis and new causal model based on more granular and precise measurements.
Several other studies [21, 41] explore and compare performance and other properties across programming languages, but do not address energy efficiency directly. These comparative studies are complementary to this work: we show here that execution time is the primary factor in energy consumption.
@@ VIII Other Related Work
Previous research also addresses the energy consumption of different data structures in certain languages, notably Java collections [47, 48, 49], or Python data manipulation libraries [50]. Data structures can be interesting candidates to study through an energy-focused lens as they can have a substantial impact on locality and memory activity, which, as Section IV-D describes, are factors in energy consumption. However, programmers often have a limited choice of data structures that fit their use case or algorithm.
Finally, past work has also investigated energy efficiency of concurrent programs, focusing on Haskell’s data sharing primitives [51] or Java’s thread management constructs [52], or analyzing the power to performance trade-offs in lock-based synchronization [53]. As Section IV-D argues, these studies confirm that the power usage patterns of the processor and the machine as a whole are more complex with parallel execution.
IX Conclusion
@@
Data supporting this paper’s evaluation is available at github.com/nicovank/Energy-Languages.